OpenStudy (kidthatbro8):

MEDAL FOR ANSWER Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (mrnood):

It tells you what f(1) is and it tells you what f(2) is It also tells you how to calculate the result when n>2 So for n = 3 it tells you f(3) = f(2) + (f(1) then f(4) = f(3) + f(2) f(5) = f(4) + f(3) SO start with f(3) since you already know f(1) and f(2) THEN you can work on to f(4) and f|(5) you then have your five terms.

OpenStudy (kidthatbro8):

it would be A?

OpenStudy (mrnood):

I think that is a guess. Please work through the calculations as I showed above: First f(3) = f(2) + (f(1) So calculate that (you know f(1) and f(2) from the question....

OpenStudy (mrnood):

And gone offline as soon as I asked for some input.... :-(

OpenStudy (kidthatbro8):

sorry ^^ it would be C?

OpenStudy (mrnood):

I am not going to confirm or deny your guesses. Have you worked out the values as indicated above? Show a little of your work and I will let you know if it's correct